A spring with a spring constant of 950 N/m is compressed 0.20 m. What speed can it give to a 1.5 kg ball when it is released?


Please answer using grade 11 Conservation of Energy and Momentum please!



1
Expert's answer
2020-05-08T16:18:38-0400

The potential energy of a compressed spring is converted into kinetic energy of the ball. We write this conservation law

"\frac{k \cdot x^2}{2}=\frac{m \cdot v^2}{2}"

then write

"v=\sqrt{\frac{k \cdot x^2}{m}}=x \cdot\sqrt{\frac{k }{m}}=0.2 \cdot\sqrt{\frac{950 }{1.5}}=5.033m/s"